# Detecting Document-level Paraphrased Machine Generated Content: Mimicking Human Writing Style and Involving Discourse Features

### v1

@@ 5 DTransformer
After preparing the structural and content features, the code sequences will be input into the decoder. The decoder is similar to a standard Transformer decoder block, with modifications made to cross-attention mechanism. We use $[CLS]$ to generate $Q$, $Code$ to generate $K,V$. The final layer of the decoder will output a probability using a linear network layer and the softmax function to determine the final classification as either MGC or HPC.
6 Implementation Details
We prepared the M4, paraLFQA, and paraWP datasets by randomly splitting each into 80% for training, 10% for validation, and 10% for testing with a fixed random seed. The Plagbench dataset was randomly divided into 60% for training, 20% for validation, and 20% for testing due to the small size of the Plagbench dataset, the standard 10% test split makes it difficult to accurately measure the model’s performance. Since the MGC is generated from HPC, we ensure that the MGC and HPC are split equally to achieve a balanced dataset. Additionally, our MhBART model requires long-form paraphrased texts paired with their original HPC counterparts, with similar meanings between the HPC and MGC. Therefore, we use the M4 dataset only for DTransformer evaluation, specifically utilizing the M4-WikiHow subset, which contains long documents.
@@ 6 Implementation Details
For MhBART and DTransformer, we initialized our model parameters using the bart-large-cnn model [^9] https://github.com/inferless/Facebook-bart-cnn and Kaiming uniform initialization He et al. (2015), respectively. The hyperparameters and optimizer selection are listed in Appendix F.
7 Evaluation
We present the classification accuracy for both our models, alongside the results from the baseline and SOTA academic and commercial approaches, namely RoBERTa Large, GPTZero, Copyleaks, and Winston AI. Given that the dataset is balanced (50% MCG and 50% HPC), accuracy is sufficient as an evaluation metric (a majority class baseline would achieve 50% accuracy). The results are shown in Table 1. Although AUC-ROC is widely used in detection tasks, it is particularly valuable for measuring performance when the cases are imbalanced while our dataset are balanced dataset. We present an AUC-ROC graph for the paraLFQA dataset as an example, demonstrating results consistent with the accuracy metric shown in 5.
Figure 5: AUC-ROC line for paraLFQA dataset of 6 models. It shows that the DTransformer has the highest AUC-ROC score.
The RoBERTa Large detector classified most cases as HPC, resulting in an accuracy around 50%. Subsequently, we discuss the results and analyze the effectiveness of our proposed models.
Table 1: Binary classification accuracy for MhBART and DTransformer compared with baseline and SOTA commercial models on both datasets.
Model paraLFQA paraWP Revised paraWP Plagbench M4-wikihow
GPTZero 0.52 0.95 0 0.69 0.57
Copyleaks 0.50 0.67 0.4 0.59 0.91
Winston ai 0.51 0.50 0.6 0.43 0.93
RoBERTa Large 0.50 0.50 0.2 0.54 0.49
MhBART (BLEU) 0.54 0.78 0.8 0.63 -
MhBART (Cosine) 0.53 0.75 - 0.72 -
DTransformer (PDTB2.0) 0.67 0.99 0.7 0.69 0.95
DTransformer (PDTB3.0) 0.70 0.99 0.7 0.61 0.94
7.1 MhBART analysis
MhBART outperforms RoBERTa Large on both document-level paraphrased datasets as it can be seen in Table 1 in bold score, and performs better on the Plagbench and paraLFQA dataset across all other detectors, achieving 0.72 and 0.54 accuracy, respectively. This indicates that MhBART can learn some implicit features, such as writing styles, in the output text. The output texts tend to have a writing tone that aligns with their related HPC versions. Examples are in Table 2. We hypothesize that the Plagbench dataset may exhibit more distinct writing style features, making it easier for MhBART to learn and thereby achieving the best performance when using cosine similarity as the evaluation metric. For the other datasets, MhBART performs slightly worse than DTransformer, as these datasets contain relatively long texts that DTransformer may be better at taking advantage of the discourse relations. Additionally, BLEU is better suited for capturing format similarity. Beyond, the complexity of these datasets may pose challenges for BERT in encoding text for the cosine similarity calculation.
Table 2: Output text examples. HPC and MGC given as inputs to MhBART and their outputs. For HPC, the output retains more content, whereas for MGC, the output tends to use fewer “ing" forms and has a more human-like writing style.
Original Texts Output Texts
HPC: The wet marble floor pressed on his cheek like a thousand hands slapping his face frozen in time. The wet marble floor pressed on his cheek like a thousand hands slapping his face frozen in time.
MGC: He woke up on the wet marble floor, feeling the discomfort of the cold surface against his cheek. He woke up on the wet marble floor. The floor was cold against his cheek, and he winced at the sensation.
For the paraWP dataset, MhBART may not be the best model, possibly due to the issue with the classification component that it cannot find the threshold of binary classification bound. There might be better classification models to find this threshold, which could provide higher detection accuracy. We provide an additional plot to illustrate the difference in BLEU1 scores between the output of MhBART and the input texts (MGC or HPC). This is shown in Appendix D, suggesting the ineffectiveness of the classification model.
Another possible reason is because paraWP has explicit features that are easily detected by GPTZero. These features include, but are not limited to, the following writing styles: a) using present or past participles as adverbial clauses and using appositives, and b) using metaphors and parallelism in sentences. To test this hypothesis, on a sample of data we removed these writing patterns by reducing the number of adverbial clauses with -ing and -ed forms, appositives, and parallel sentences. This simplifies sentences by using essential logical connectors (such as so, but, because, and) and by deleting repeated elements. We did that on 10 paragraphs (3 are long documents), and it turns out that the GPTZero cannot classify the revised version correctly but MhBART would not be influenced negatively using BLEU-1/4 as a similarity metric. One of the revised example comparisons is shown in Table 6. It shows that the MhBART learns more than simple explicit features which GPTZero might rely on for detection.
We think another factor contributing to the performance of commercial tools is the dataset-specific bias in their outputs. Evidence for this is seen in the lack of a single dominating commercial tool across all of datasets. The exceptional performance of certain tools may be attributed to the specific training samples they were exposed to. For instance, GPTZero might have been trained on a larger number of samples generated by ChatGPT using prompts similar to those used to produce paraWP, but fewer for M4. Consequently, the model’s performance may not necessarily reflect an effective feature extraction strategy, but rather the effectiveness of its training data with specific types of prompts.
7.2 DTransformer analysis
The results presented in the Table 1 demonstrate that the DTransformer outperforms all other models across the paraLFQA, showing its efficacy in extracting features from document-level texts. ParaLFQA is discourse paraphrased so DTransformer could take advantage of discourse inforation while other models fail to pay attantion to. For the Plagbench dataset, it performs slightly worse than MhBART, likely because the dataset lacks a sufficient number of distinct discourse features based on the prompts used for data generation. These performances highlight the significance of incorporating discourse features. The critical of discourse features is further validated through comparative experimentation using PDTB 2.0 and 3.0, where PDTB 3.0 consistently outperforms 2.0. This superiority is attributable to PDTB 3.0’s more comprehensive discourse annotations, which better represent discourse features effectively. Specifically, the best-performing DTransformer model utilizing PDTB 3.0 shows a 4%, 15.5% 1.5% absolute improvement over existing commercial models on the paraWP, paraLFQA, and M4 datasets, respectively.
Our result reveals that DTransformer exhibits high performance on the paraWP dataset, achieving an accuracy of 99%. This indicates that DTransformer possesses a remarkable capability for detecting MGC, even within the constraints of a challenging dataset. Furthermore, the design of the generation prompt plays a crucial role. Specifically, we ensured that the generated MGCs were predominantly distinct from their corresponding HPCs at the discourse feature level, while controlling for other variables such as semantic meaning and word count. This alignment with DTransformer’s capabilities accounts for the exceptionally high accuracy observed.
We also performed a case study, shown in Appendix E to assess the effectiveness of the PDTB.
In addition, we conducted an ablation study focusing on discourse features. To successfully extract document-level semantic meaning as DTransformer does, we trained the original HiBERT model Zhang et al. (2019) on the paraWP dataset. However, the predicted accuracy with the hierarchical architecture model alone is 0.53 and 0.49
for the PDTB3.0 and PDTB2.0 annotated datasets, respectively on paraWP. This finding highlights the importance of incorporating both discourse and semantic features for effective document-level MGC detection. Additionally, it highlights that the accuracy gain for DTransformer is primarily attributable to discourse features.
In light of the models’ high performance, we aim to evaluate the scalability of our models. We conducted five independent experiments for each model, utilizing different random seeds. DTransformer with PDTB 2.0 resulted in a mean test accuracy of 0.67 and a standard deviation of 0.03 on the paraLFQA dataset. The DTransformer with PDTB 3.0 produced a mean test accuracy of 0.70 and a standard deviation of 0.02. These results indicate the high mean accuracy and consistency of both models, demonstrating good scalability. The lower standard deviation of PDTB 3.0 suggests more stable performance across different runs.
8 Conclusions
In this work, we introduced two new datasets aiming to provide insights into the challenges in current work aimed at detecting MGC, particularly for document-level paraphrased machine-generated texts: paraLFQA, emphasizing the discourse difference, and paraWP, a long dataset with highly similar between HPC and MGC. Further, we introduced two architectures, MhBART and DTransformer, outperforming current commercial and non-commercial models. The former incorporates human-written styles, the latter utilizes discourse features. We reported that our models outperform the competition across the datasets paraLFQA, paraWP, Plagbench, and M4. Additionally, we identified patterns in writing styles that affect the performance of existing detectors and show that our detectors are more robust when faced with these patterns. Future work could focus on integrating both models into a end-to-end detection system, enhancing efficiency in detection workflows.
Limitations
A. We only report three sentence patterns, or writing styles, that affect GPTZero’s detection. It could be a future direction to find ways to embed these writing styles into automatic detection methods.
B. We excluded non-latin alphabet languages, such as Tibetan, as the current focus is on certain behaviour of models in a simple testbed without complications. The intention is to avoid complications arising from challenges associated with non-latin alphabet languages. During our study, every text with non-latin characters was removed from the dataset.
C. We eliminated obvious differences between HPC and MGC regarding writing habits. This includes instances of consecutive emotional markers like “!!!!!!” or artificial modifications of words like “su–per!”, used merely to express a pause. We detect if there are invalid markers within the words, then, we removed this sample, as the goal of the study was to specifically focus on learning intrinsic features, such as text structure and writing style.
D. We attempted to embed methods for automatically removing sentence patterns that fool GPTZero, such as using GPT-3.5 and GPT4/4o to rephrase the entire corpus. However, any prompts we tried such as “Rewrite the paragraph to replace all the appositive and participle clauses including ing or ed. Do not change, add or delete content words. Retain other parts of the language expression.” did not remove all the present or past participle clauses. This might be as certain prompts cannot apply to all texts.
E. Due to GPU limitations, we selected a manageable batch size and maximum sentence length for our dataset. As a result, our model has been trained on the target dataset but would benefit from additional training for broader, out-of-domain commercial applications. While commercial models have likely not been exposed to the specific datasets we investigated, they leverage significantly larger datasets in their training processes, though the specifics of this data are not publicly disclosed.
Ethics Statement
Our work has important ethical implications. The advancements in MGC detection not only help to foster a more honest academic environment but also align with the ethical responsibility to maintain integrity in education. As conscious testing for MGC shows initial success, continuing this trend supports academic honesty and fairness for all students. Moreover, effective MGC detection plays a crucial ethical role in combating the spread of fake news and misleading information, which promotes truth and trust in public discourse. In terms of the innovation of music and drawings, detecting MGCs helps preserve the authenticity of artistic expression, preventing forgery and ensuring rightful appreciation to creativity. Through these efforts, we aim to uphold ethical standards across various fields and mitigate the misuse of MGC.
Appendix A Writing prompts data linguistic language features
We investigate basic linguistic features of paraWP: most commonly used words, sentence lengths, lexical diversity, token-type ratio, and further more. The graphs below indicate the top 20 frequently used terms in the HPC (Fig. 6) and MGC (Fig.  7) datasets. It could be seen that the most frequently used words are similar and the number of times the words used is also similar. There is not much difference in these feature.
Figure 6: Top 20 most used word frequencies calculated in the HPC corpus.
Figure 7: Top 20 most used word word frequencies calculated in the MGC corpus. It is used to compare the diversity of words between MGC and HPC.
Additionally, other linguistic features are presented in Tables 3, 4, and 5.
Data Avg. Sent. Length Lexicon Diversity
MGC 486.34 0.498
HPC 559.44 0.474
Table 3: Average Sentence Length and Lexicon Diversity. We calculate the average number of words in one document (Average Sentence Length) and the proportion of different words in total words (Lexicon Diversity). These are two other linguistic features for MGC and HPC linguistic features comparison.
Data Avg. Sent. ing PTCP ed PTCP CC
MGC 28.04 13.57 10.84 16.14
HPC 41.75 11.75 9.93 17.84
Table 4: Average number of sentences, present & past participle and connectives in the test set. We calculate the average number of sentences in one document (Avg. Sent.) and also for present, past participles, and connectives given their total number of characters are similar. These are the comparisons of the sentence patterns between HPC and MGC.
The POS tag ratios provide syntactic information, and the average lexicon diversity is calculated by dividing the number of distinct words used by the total number of words. There are no significant differences in any aspect. Overall, explicit features such as syntactic or content features of MGCs and HPCs are similar. It is challenging to detect MGC or HPC relying only on these.
POS MGC Ratio HPC Ratio
Det 0.1099 0.0993
Noun 0.2424 0.2209
Adj 0.0613 0.0674
Verb 0.2129 0.2235
Adp 0.1156 0.1045
Pron 0.1217 0.1346
Num 0.0051 0.0079
Conj 0.0379 0.0349
Prt 0.0394 0.0356
Adv 0.0531 0.0698
X 0.0007 0.0018
Table 5: POS Tag Ratios for MGC and HPC. We use universial POS tags for word classification and calculate their respective proportions. This serves as an indicator of superficial syntactic features.
Appendix B Universial authorship representation (UAR) representation experiments
The UAR is utilized as implicit features of writing style, proposed as stylistic representation. We calculate the difference score of cosine similarity based on UAR, and the results are quite poor, with accuracy scores of 0.623, 0.574, and 0.518 for paraWP, Plagbench, and paraLFQA, respectively. We visualize the UAR using UMAP, following the methodology proposed in the UAR work Soto et al. (2021), as shown in Fig. 8.
Figure 8: UMAP visualization of UAR representations for four types of texts: HPC (top-left), intermediate output with HPC and MGC as input (top-right), and MGC (bottom-left) in paraLFQA (up) and paraWP (down). Each plot shows approximately 300 texts for each type of texts.
This figure provides insight into the suboptimal performance of MhBART when using UAR. In paraLFQA, there is a lack of distinct separation among the four text types, whereas in paraWP, the intermediate outputs exhibit a higher degree of similarity. This phenomenon is attributed to the limitations of UAR in effectively capturing the entire writing style and its challenges in fine-tuning for our specific task. Future work should focus on developing an automated method to evaluate this implicit feature.
Appendix C Revised MGC example
The comparison can be seen in Table 6. Changes are marked in font in the texts below. Long texts without changes are indicated by . If the AI percentage surpasses 50%, GPTZero should classify the content as MGC. Conversely, if it is 50% or lower, GPTZero should classify it as HPC.
Table 6: Revised MGC example. The font parts indicate all the manual revisions. The entire text is around 450 words, and we show all changes within significant contexts. Parts that are not shown have not been changed.
Original MGC: 100% AI detected by GPTZero Revised MGC: 44% AI detected by GPTZero
He woke up on the wet marble floor, feeling the discomfort of the cold surface against his cheek. Ignoring the rain outside, he struggled to gather his thoughts, but the memories remained veiled by a black mist. Realizing he didn’t know where he was, he searched the room and found a bottle of vodka. Taking a few swigs, he used it as a distraction from his forgotten past. With nothing else to do, he looked at himself in the mirror. His appearance suggested he could be a successful mattress salesman or a lucky gambling bum. His face was unshaven, his stomach round yet firm, and his hair long and unruly. However, his eyes still held a hint of intensity, like a rusted dagger that could still cut. In frustration, he cursed the name "DiCaprio," resenting the fact that despite his efforts to bury himself in alcohol and drugs, he was still associated with the actor. Even after bankruptcy, the royalties from Titanic and other projects kept pouring in, making it impossible for him to escape his past. He threw the bottle at the mirror, but it slipped from his wet hands and only made a thud against the wall. Defeated, he slumped against the floor and noticed the rain entering the room. Frustrated and confused, he wondered what had happened the night before. As he sat there, the pieces of his fragmented memory started to come together. He had been involved in a violent altercation, and the consequences were now staring back at him through the broken window. Fear and regret washed over him as he tried to piece together the events leading up to this moment. The pain in his body matched the pain in his mind, and he knew he couldn’t escape the consequences of his actions. The room, once a place of mystery, now became a haunting reminder of the darkness within him. He slowly stood up, his body aching with every movement. Determined to uncover the truth, he stepped over the broken glass and ventured out into the rain-soaked streets. The journey to unravel the mysteries of that fateful night had just begun, and he knew he had to confront his past, no matter how painful it might be. He woke up on the wet marble floor, and felt the discomfort of the cold surface against his cheek. He ignored the rain outside. He struggled to gather his thoughts, but the memories remained veiled by a black mist. Because he realized he didn’t know where he was, he searched the room and found a bottle of vodka. He took a few swigs, and used it as a distraction from his forgotten past. He looked at himself in the mirror but had nothing else to do. His appearance suggested he could be a successful mattress salesman or a lucky gambling bum. His face was unshaven. His stomach was round yet firm. His hair was long and unruly. However, his eyes still held a hint of intensity, like a rusted dagger that could still cut. He was in frustration, and cursed the name "DiCaprio." This resents that despite his efforts to bury himself in alcohol and drugs, he was still associated with the actor. Even after bankruptcy, the royalties from Titanic and other projects kept pouring in. This made it impossible for him to escape his past. He threw the bottle at the mirror, but it slipped from his wet hands and only made a thud against the wall. He was defeated so he slumped against the floor. He noticed the rain entering the room. He was frustrated and confused. He wondered what had happened the night before. He sat there. The pieces of his fragmented memory started to come together. He had been involved in a violent altercation, so the consequences were now staring back at him through the broken window. Fear and regret washed over him when he tried to piece together the events. The pain in his body matched the pain in his mind, but he knew he couldn’t escape the consequences of his actions. The room is a place of mystery, but now became a haunting reminder of the darkness within him. He slowly stood up. His body aches with every movement. He was determined to uncover the truth, so he stepped over the broken glass and ventured out into the rain-soaked streets. The journey to unravel the mysteries of that fateful night had just begun. He knew he had to confront his past, no matter how painful it might be.
Appendix D BLEU plots for MhBART results
This appendix provides additional plots to show the difference score (BLEU1) between the output of MhBART and input (MGC or HPC texts) shown in Fig. 9
Figure 9: The BLEU score as a difference score for 128 cases in the validation set of WritingPrompts shows the differences in divergence between HPCs and MGCs with the outputs of MhBART.
This evidence suggests that the model learns features from HPCs because the output tends to mimic and resemble HPCs, given that the overall bleu line trend of HPCs is higher than the red line of MGCs, even though there is no clear threshold. There might be better classification models to find this threshold, which could provide higher detection accuracy.
Appendix E Case study for the DTransformer result
We performed a case study, as shown in Table 7, to assess the effectiveness of the PDTB. For this evaluation, we employed the DTransformer with PDTB 3.0 as the representative model. The case study analyzed a paragraph from the paraLFQA dataset, originally labeled as MGC. This paragraph was misclassified by GPTZero and MhBART as HPC, but it was correctly classified as HPC by DTransformer.
Table 7: DTransformer test example: A paragraph from paraLFQA is selected and its corresponding PDTB code prediction is presented. DTransformer made correct classifications while MhBART and GPTZero classified falsely.
Text PDTB code
Basically, they are related to the amount of “advantage" which a lever and a winch give you. [CLS]
The PDTB codes are intuitively accurate. For instance, the ‘expansion.level-of-detail’ code corresponds to the elaboration on the ‘lever’ in the second sentence, following the ‘lever’ discussed in the previous sentence. Additionally, the ‘contingency.cause’ code is marked by sentences containing conjunctions such as ‘but’ and ‘so’ to indicate cause and effect relationships. The ‘expansion.instantiation’ code is identified by the presence of the phrase ‘for example.’ These examples demonstrate that PDTB coding aligns well with human judgment, providing evidence for the effectiveness of discourse features in detecting MGC. Conversely, we hypothesize that MhBART and GPTZero produced incorrect classifications due to their relative emphasis on writing style and other semantic features rather than on discourse features.
Appendix F Hyperparameter and opmimizer for our models
Table 8: Hyperparameters, tokenizer and optimiser for MhBART training.
Hyperparameter MhBART
max-sentence-length 1024
batch 1
learning rate 2e-5
epochs 10
early stop epochs 3
optimizer AdamW
tokenizer BPE
@@ Appendix F Hyperparameter and opmimizer for our models
Table 9: Hyperparameters, tokenizer and optimizer for DTransformer.
@@
The batch sizes differ between the two datasets of DTransformer training is due to GPU memory limitations. We use a batch size of 4 for paraWP and 16 for paraLFQA, reflecting the varying dataset sizes, with paraLFQA being slightly shorter in length. The device we use is a Tesla V100-PCIE-32GB. Also, we could only use batch size 1 for the MhBART experiments due to the GPU memory limitations. This constraint likely contributes to the suboptimal performance of our MhBART model, as batch size plays a crucial role in model training.

### v2

@@ IV Model: DTransformer
After preparing the structural and content features, the code sequences will be input into the decoder. The decoder is similar to a standard Transformer decoder block, with modifications made to the cross-attention mechanism. We use $[CLS]$ to generate $Q$, CodeCode to generate $K,V$. The final layer of the decoder will output a probability using a linear network layer and the softmax function to determine the final classification as either MGC or HPC.
V Implementation Details
We prepared the M4, paraLFQA, and paraWP datasets by randomly splitting each into 80 % for training, 10 % for validation, and 10% for testing with a fixed random seed. The Plagbench dataset was randomly divided into 60 % for training, 20 % for validation, and 20 % for testing due to the small size of the Plagbench dataset, the standard 10 % test split makes it difficult to accurately measure the model’s performance. Since the MGC is generated from HPC, we ensure that the MGC and HPC are split equally to achieve a balanced dataset.
@@ V Implementation Details
For DTransformer, we initialized our model parameters using the bart-large-cnn model [^10] https://github.com/inferless/Facebook-bart-cnn and Kaiming uniform initialization [54], respectively. The hyperparameters and optimizer selection are listed in Table II.
@@
TABLE II: Hyperparameters, tokenizer and optimizer for DTransformer.
@@
The batch sizes differ between the two datasets of DTransformer training is due to GPU memory limitations. We use a batch size of 4 for paraWP and 16 for paraLFQA, reflecting the varying dataset sizes, with paraLFQA being slightly shorter in length. The device we use is a Tesla V100-PCIE-32GB.
VI Evaluation
We present the classification accuracy for both our models, alongside the results from the baseline and SOTA academic and commercial approaches, namely RoBERTa Large, GPTZero, Copyleaks, and Winston AI. Given that the dataset is balanced (50% MCG and 50% HPC), accuracy is sufficient as an evaluation metric (a majority class baseline would achieve 50% accuracy). The results are shown in Table III. The RoBERTa Large detector classified most cases as HPC, resulting in an accuracy around 50%. Subsequently, we discuss the results and analyze the effectiveness of our proposed models.
TABLE III: Binary classification accuracy for DTransformer compared with baseline and SOTA commercial models on both datasets.
Model paraLFQA paraWP Revised paraWP Plagbench M4-wikihow
GPTZero 0.52 0.95 0 0.69 0.57
Copyleaks 0.50 0.67 0.4 0.59 0.91
Winston ai 0.51 0.50 0.6 0.43 0.93
RoBERTa Large 0.50 0.50 0.2 0.54 0.49
DTransformer (PDTB2.0) 0.67 0.99 0.7 0.69 0.95
DTransformer (PDTB3.0) 0.70 0.99 0.7 0.61 0.94
The results presented in Table III demonstrate that the DTransformer outperforms all other models across datasets, especially the paraLFQA, showing its efficacy in extracting features from document-level texts. For ParaLFQA, the discourse paraphrasing is particularly high, such that DTransformer could take advantage of discourse information while other models fail to pay attention to this aspect. These performances highlight the significance of incorporating discourse features. The criticality of discourse features is further validated through comparative experimentation using PDTB 2.0 and 3.0, where PDTB 3.0 consistently outperforms 2.0. This superiority is attributable to PDTB 3.0’s more comprehensive discourse annotations, which better represent discourse features effectively. Specifically, the best-performing DTransformer model utilizing PDTB 3.0 shows a 4 %, 15.5 % 1.5 % absolute improvement over existing SOTA models on the paraWP, paraLFQA, and M4 datasets, respectively, which is strong evidence for DTransformer’s ability for general document-level detection.
Our results reveal that DTransformer exhibits high performance on the paraWP dataset, achieving an accuracy of 99 %. This indicates that DTransformer possesses a remarkable capability for detecting MGC, even within the constraints of a challenging dataset. Furthermore, the design of the generation prompt plays a crucial role. Specifically, we ensured that the generated MGCs were predominantly distinct from their corresponding HPCs at the discourse feature level, while controlling for other variables such as semantic meaning and word count. This alignment with DTransformer’s capabilities accounts for the exceptionally high accuracy observed.
We also performed a case study, shown in Table IV to assess the effectiveness of the PDTB. For this evaluation, we employed the DTransformer with PDTB 3.0 as the representative model. The case study analyzed a paragraph from the paraLFQA dataset, originally labeled as MGC. This paragraph was misclassified by GPTZero as HPC, but it was correctly classified as HPC by DTransformer.
TABLE IV: DTransformer test example: A MGC-paragraph from paraLFQA is selected and its corresponding PDTB code prediction is presented. DTransformer made correct classifications while GPTZero classified falsely.
Text PDTB code
Basically, they are related to the amount of “advantage” which a lever and a winch give you. [CLS]
The PDTB codes are intuitively accurate. For instance, the ‘expansion.level-of-detail’ code corresponds to the elaboration on the ‘lever’ in the second sentence, following the ‘lever’ discussed in the previous sentence. Additionally, the ‘contingency.cause’ code is marked by sentences containing conjunctions such as ‘but’ and ‘so’ to indicate cause and effect relationships. The ‘expansion.instantiation’ code is identified by the presence of the phrase ‘for example.’ These examples demonstrate that PDTB coding aligns well with human judgment, providing evidence for the effectiveness of discourse features in detecting MGC. Conversely, we hypothesize that GPTZero produced incorrect classifications due to its relative emphasis on writing style and other semantic features rather than on discourse features.
In addition, we conducted an ablation study focusing on discourse features. To successfully extract document-level semantic meaning as DTransformer does, we trained the original HiBERT model [36] on the paraWP dataset. However, the predicted accuracy with the hierarchical architecture model alone is 0.53 and 0.49 for the PDTB3.0 and PDTB2.0 annotated datasets, respectively on paraWP. This finding highlights the importance of incorporating both discourse and semantic features for effective document-level MGC detection. Additionally, it highlights that the accuracy gain for DTransformer is primarily attributable to discourse features.
In light of the models’ high performance, we aim to evaluate the scalability of our models. We conducted five independent experiments for each model, utilizing different random seeds. DTransformer with PDTB 2.0 resulted in a mean test accuracy of 0.67 and a standard deviation of 0.03 on the paraLFQA dataset. The DTransformer with PDTB 3.0 produced a mean test accuracy of 0.70 and a standard deviation of 0.02. These results indicate the high mean accuracy and consistency of both models, demonstrating good scalability. The lower standard deviation of PDTB 3.0 suggests more stable performance across different runs. Additionally, the high accuracy of the revised paraWP serves as evidence that DTransformer is highly robust against surface-level attacks.
Based on this analysis, we argue that our findings are theoretically supported: machines struggle to maintain a stable discourse structure in document-level text generation, whereas humans can preserve a clear central theme when writing long texts. This is consistent with previous research suggesting that LLMs cannot effectively process and comprehend long documents. Instead, they tend to segment texts into chunks and process them sequentially, leading to a loss of contextual coherence [55, 56]. Our approach is effective because we explicitly provide this structural information to the model.
VII Conclusions
In this work, we introduced two new datasets aiming to provide insights into the challenges in current work aimed at detecting MGC, particularly for document-level paraphrased machine-generated texts: paraLFQA, emphasizing the discourse difference, and paraWP, a long dataset with high similarity between HPC and MGC. Further, we introduced a novel architecture, DTransformer, which utilises discourse features, outperforming current commercial and non-commercial models across the datasets paraLFQA, paraWP, Plagbench, and M4. Future work could focus on training on more datasets for out of domain tests, enhancing robustness of the model.
Appendix A Writing prompts data linguistic language features
We investigate basic linguistic features of paraWP: most commonly used words, sentence lengths, lexical diversity, token-type ratio, and further more.
We analyzed the top 10 most frequently used terms in the HPC (e.g., like, time, back, know, said, man, see, eyes, get, around) and MGC (time, life, man, world, like, back, eyes, face, help, find in descending order of frequency) datasets. The results indicate that the most frequently used words are similar across both datasets, and their frequency of occurrence is also comparable. Overall, there is little variation in this feature.
The MGC and HPC datasets have average sentence lengths of 486.34 and 559.44, respectively, and lexical diversity (measured as the proportion of unique words to total words) of 0.498 and 0.474, indicating similar characteristics. Additionally, other linguistic features are presented in Tables V and VI, showing similar features and challenges for detecting them.
Data Avg. Sent. ing PTCP ed PTCP $C$
MGC 28.04 13.57 10.84 16.14
HPC 41.75 11.75 9.93 17.84
TABLE V: Average number of sentences, present & past participle and connectives in the test set.
The POS tag ratios provide syntactic information, and the average lexicon diversity is calculated by dividing the number of distinct words used by the total number of words. There are no big differences in any aspect. Overall, explicit linguistic features of MGCs and HPCs are similar. It is challenging to detect MGC or HPC relying only on these.
POS MGC Ratio HPC Ratio
Det 0.1099 0.0993
Noun 0.2424 0.2209
Adj 0.0613 0.0674
Verb 0.2129 0.2235
Adp 0.1156 0.1045
Pron 0.1217 0.1346
Num 0.0051 0.0079
Conj 0.0379 0.0349
Prt 0.0394 0.0356
Adv 0.0531 0.0698
X 0.0007 0.0018
TABLE VI: POS Tag Ratios for MGC and HPC. We use universial POS tags for word classification and calculate their respective proportions. This serves as an indicator of superficial syntactic features.